# A Dirichlet stochastic block model for composition-weighted networks

### v1

@@ 5.4 Model selection performance
We assess the model selection performance using the integrated completed likelihood (ICL) detailed in Section 3.4. Table 1 reports the confusion matrices of the true number of clusters $K$ and the optimal number of clusters $\hat{K}$ as selected by ICL.
Table 1: Choice of the number of clusters using integrated completed likelihood (ICL). Results indicate the number of times each number of clusters has been selected by the criterion, based on 50 artificial networks.
@@
We observe that ICL is successful in selecting the correct number of clusters in larger networks with any level of parameter homogeneity, which is to be expected provided that both the latent structure recovery and the parameter estimation improve with network size. It is also not surprising that, in smaller networks with higher levels of homogeneity and/or higher numbers of clusters, the model selection becomes a much more complex task, and the number of clusters tends to be underestimated by the ICL. This is due to the fact that in cases with medium and high levels of homogeneity the clusters are much harder to separate, coupled with the difficulty of accurately estimating the parameters in smaller networks, as seen in Sections 5.3 and 5.2 respectively.
@@ 6.1 Erasmus programme data
The Erasmus programme is the most popular university student exchange programme in Europe, spanning thousands of Higher Education institutions in more than 30 countries. Gadár et al., (2020) have produced an extensive data set of student, staff and teacher mobility in the Erasmus network by combining data from a number of different sources.
The aim of our analysis is to understand global mobility patterns in the Erasmus network, focusing on student exchanges. Due to significant differences in the student populations of the European countries, analysis of raw counts of students per pair of countries may not provide particular insights regarding the popularity of destinations, as the biggest countries dominate simply due to their overall hosting capacity.
After aggregating the student numbers to country level for the 2012-2013 academic year, a directed network with countries as nodes and proportions of students as edge weights is constructed, with the proportions summing to 1 for the sending countries. The Erasmus network is very dense ($\approx 87\%$ of all possible edges are present), and there is no reason to believe that the remaining edges are not allowed to exist since all countries in the Erasmus network are allowed to exchange students. Therefore, instead of treating the edges as truly missing, we assign a value of $0.001$ to the interactions with zero count and treat the network as fully connected. In compositional data analysis, zero values present a great challenge and are an active area of research since standard tools are not designed to handle zero-valued compositional parts (Filzmoser et al.,, 2018). Replacing zeros by some small value is a standard approach in the case of zero counts that works generally well, especially considering its simplicity. Since in our case $0.001$ is added to the original count data rather than to the compositions themselves, the final results should be robust relative to the magnitude of the added constant.
@@
Figure 4(a) illustrates the solution with 3 clusters on a map of Europe, noting the membership of each cluster on the side as smaller countries are difficult to see on the plot. Figure 4(b) represents the total percentage flows between the clusters, based on the estimated matrix $\mathbf{\hat{V}}$. The estimate of the parameter matrix $\mathbf{A}$ as well as the expected node-to-node and cluster-to-cluster exchange matrices $\mathbf{\hat{W}}$ and $\mathbf{\hat{V}}$ from Equations (5) and (6) respectively are provided in Appendix E.
Figure 4: Clustering solution for DirSBM with 3 clusters: 4(a) map of Europe with countries coloured by cluster assignment; 4(b) chord diagram of total percentage flows between clusters of countries based on $\mathbf{\hat{V}}$. On the left, S1, S2 and S3 denote the sending cluster (1, 2 and 3). Percentages departing from the sender sum up to 100. On the right, R1, R2 and R3 denote the receiving cluster (1, 2 and 3).
@@
Figure 5: Clustering solution for DirSBM with 5 clusters: 5(a) map of Europe with countries coloured by cluster assignment; 5(b) chord diagram of total percentage flows between clusters of countries based on $\mathbf{\hat{V}}$. On the left, S1-S5 denote the sending cluster (numbered 1 to 5). Percentages departing from the sender sum up to 100. On the right, R1-R5 denote the receiving cluster (numbered 1 to 5).
@@
The second best solution according to ICL with $K=5$ is presented in Figure 5, and the parameter estimates as well as matrices $\mathbf{\hat{V}}$ and $\mathbf{\hat{W}}$ are available in Appendix E. The red cluster, now consisting of only Germany, Spain and France, is the cluster of countries receiving the highest percentages of students from the countries of all clusters, ranging between the expected 8.1% to 19.6% per country pair, and the second largest total shares, dominated by the blue cluster by a small margin. It should be noted that the total shares received by the red cluster are only split between 3 countries instead of 11 like in the case of the blue cluster. This cluster is also characterised by retaining 39.1% of its students, and sending 43.3% to the blue cluster, implying a lack of diversity of exchange destinations among its students. The blue cluster can be described as containing reasonably popular exchange destinations, but not quite as popular as the red countries, and, similarly to the red cluster, the majority of students go on exchange within the cluster itself (42.2%) or to the red cluster (38.1%). The green cluster, despite consisting of 9 countries (8 Eastern European countries and Cyprus), does not attract large shares of students (total cluster exchange shares are between 2.4% and 10.6%), which may be due to the fact that such countries are less well known exchange destinations. Green countries have a somewhat more diverse preference profile in comparison to the first two clusters, as the parameter estimates are much closer to each other, but there is still a strong preference for the red and blue cluster countries. Purple countries, which are Switzerland, Hungary, Ireland, Norway and Turkey, can be described as medium-preference exchange destinations. They have a strong preference for the red and blue cluster countries, and very limited interest in countries within its own cluster, as well as in green and orange clusters. The orange cluster is the one having the most homogeneous preferences, as indicated by the smallest differences between the parameter estimates, and its countries tend to be among the least popular destinations for exchange. This can possibly be explained by the fact that these countries are either very small, are located very far from the centre of Europe or are not members of the European Union (as of the 2012-2013 academic year). The green, purple and orange clusters retain very small shares or transfers, 14.2%, 2.4% and 5.6% respectively, while the majority of students from these clusters tend to go on exchange to countries in the red and blue clusters. A final observation is that there seems to be a connection between the exchange preferences of Erasmus students and the level of economic development of the hosting country or the prestige of going on exchange to specific destinations. The countries with stronger economies tend to be grouped together and to exhibit higher expected country-to-country exchange proportions, whereas countries with weaker economies or newer members of the programme, such as Turkey and Croatia, tend to be separated from the more established members.
@@ 6.2 London bike sharing data
We illustrate DirSBM in application to the London bike sharing data from 2014 provided by the Transport for London, (2016). We start by counting the number of trips taken between pairs of stations and identifying the busiest stations for our analysis. To do so, we look at the top 100 start and end stations in terms of volume of bikes exchanged and consider their intersection, i.e. the stations that are both among the most popular start and end points of the journey, resulting in an overall network of 85 stations. The trips only between these 85 stations account for over 10% of the total year’s trips. The network is very dense, with just under 2% of edges having zero weights, hence, similarly to the Erasmus programme network from Section 6.1, we set the weights of zero-weighted edges equal to $0.001$. We then compute the compositional edge weights by dividing the count weights by the total number of trips taken from the start station.
We fit the DirSBMs with number of clusters between 1 and 8 clusters and select the optimal number of clusters based on the ICL. According to the criterion, the solution with $K=4$ is optimal (ICL$=27866$), followed by the solution with $K=3$ (ICL$=27608$). Figure 6 illustrates the solution with 4 clusters on a map of central London (an interactive version of the map is available here) as well as the chord diagram for total exchange proportions between clusters. The Dirichlet parameter matrix estimate $\mathbf{\hat{A}}$ as well as matrices $\mathbf{\hat{V}}$ and $\mathbf{\hat{W}}$ are provided in Appendix F.
@@
Figure 6: Clustering solution for DirSBM with 4 clusters, as selected by ICL, on London bike sharing data: 6(a) map of Central London with bike stations coloured by cluster assignment. Image created using Google My Maps software, with the interactive map available here; 6(b) chord diagram of total percentage flows between clusters of stations based on $\mathbf{\hat{V}}$. On the left, S1-S4 denote the sending cluster (numbered 1 to 4). Percentages departing from the sender sum up to 100. On the right, R1-R4 denote the receiving cluster (numbered 1 to 4).
@@ 7 Discussion
In this paper, we have introduced DirSBM, an extension of the stochastic block model to directed weighted networks with compositional edge weights that is based on direct modelling of compositional weights vectors using the Dirichlet distribution. To the best of our knowledge, no counterparts currently exist in the literature, making the proposed clustering methodology the first of its kind. We have developed an inferential procedure based on a variant of the classification expectation-maximisation algorithm for hybrid likelihood. Model selection is addressed using the integrated completed likelihood criterion. We have also proposed an alternative framework for clustering for composition-weighted networks via a log-ratio transformation of the data and subsequent application of the existing weighted SBM to the transformed data. In the simulation studies, we have shown the effectiveness of DirSBM in terms of parameter estimation and recovery of the clustering structure in a variety of settings, and have explored the use of ICL for model selection, which has proven to work reasonably well. The code implementing the modelling framework in R can be found on GitHub.
The modelling approach can be extended in multiple directions. One important extension to the DirSBM is introducing structural zeros, to enable the ability to model the absence of guaranteed edges in the network. One could use the idea from Tsagris and Stewart, (2018) that concerns Dirichlet regression with zeros, where indicators are used to “discard” zero entries of compositional vectors and subsequently fit the model to the non-zero subset of compositional data. As well as making the model more realistic and applicable to a wider range of real world data, such an extension could lead to numerically more stable performance, since the dimension of the Dirichlet observations would not scale with the number of nodes.
@@
Other developments could include investigating more educated and efficient initialisation strategies for the proposed algorithm. Although random initialisation allows to explore the solution space rather well by considering the most diverse set of starting points, thus reducing the risk of convergence to a local maximum, it does so at a computational cost. This cost could potentially be reduced if the algorithm was to be run from fewer, or ideally a single, more informed initial clustering allocation set. Other distributions for compositional data could be considered to model the sets of proportional edge weights in the network, such as the generalised Dirichlet distribution, which would be less restrictive on the compositional variance (Connor and Mosimann,, 1969).

### v2

@@ 5.4 Model selection performance
Table 1: Choice of number of clusters using integrated completed likelihood (ICL). Results indicate the number of times each number of clusters has been selected by the criterion, based on 50 artificial networks.
@@
We assess the model selection performance using the integrated completed likelihood (ICL) detailed in Section 3.4. Table 1 reports the confusion matrices of the true number of clusters $K$ and the optimal number of clusters $\hat{K}$ as selected by ICL. We observe that ICL is successful in selecting the correct number of clusters in larger networks with any level of parameter homogeneity, which is to be expected provided that both the latent structure recovery and the parameter estimation improve with network size. It is also not surprising that, in smaller networks with higher levels of homogeneity and/or higher numbers of clusters, the model selection becomes a much more complex task, and the number of clusters tends to be underestimated by the ICL. This is due to the fact that in cases with medium and high levels of homogeneity the clusters are much harder to separate, coupled with the difficulty of accurately estimating the parameters in smaller networks, as seen in Section 5.3 and 5.2 respectively.
5.5 Effect of zero-weighted edges
In this section we examine the effect of the presence of zero-weighted edges in the network on parameter estimation and clustering performance of DirSBM. To generate synthetic networks with compositional weights arising from weighted networks having a certain proportion $p_{0}$ of zero-weighted edges and a block structure, we consider the following data generating process:
  1. 1.
    Given a $K$-dimensional vector of cluster membership proportions $\boldsymbol{\theta}$, generate binary cluster allocations $\mathbf{z}_{i}\sim Multinom(1,\boldsymbol{\theta})$, for $i=1,\ldots,n$. The entries in these vectors are $z_{ik}=1$ when node $i$ is assigned to cluster $C_{k}$ and 0 otherwise. Denote with $\mathbf{Z}$ an $n\times K$ matrix of binary cluster allocations of all nodes in the network, and with $\mathbf{Z}_{-i}$ the $(n-1)\times K$-dimensional matrix of cluster allocations of all nodes in the network that are not $i$.
  2. 2.
    Let $\mathbf{A}=\{\alpha_{kh}\}_{k,h=1}^{K}$ be a $K\times K$ matrix with strictly positive values. Given the cluster allocations, generate independent gamma realisations corresponding to the raw data
    $y_{ij}\lvert z_{ik}=1,\leavevmode\nobreak\ z_{jh}=1\sim Gamma(\alpha_{kh},1)$
    for $i=1,\ldots,n$ and $j=1,\ldots,i-1,i+1,\ldots,n$.
  3. 3.
    Set a proportion $p_{0}$ of edges $y_{ij}$ equal to zero. Consequently, the resulting weighted adjacency matrix $\mathbf{Y}$ will contain a certain percentage of zero-weighted edges.
  4. 4.
    Set the zero-weighted edges equal to some small quantity $\epsilon$. Here we consider $\epsilon=0.001$.
  5. 5.
    From the resulting raw data $\mathbf{Y}$, construct compositions
    $\mathbf{x}_{i}^{*}=\bigg{(}\frac{y_{i1}}{\sum_{j\neq i}^{n}y_{ij}},\frac{y_{i2}}{\sum_{j\neq i}^{n}y_{ij}},\ldots,\frac{y_{i(i-1)}}{\sum_{j\neq i}^{n}y_{ij}},\frac{y_{i(i+1)}}{\sum_{j\neq i}^{n}y_{ij}},\ldots,\frac{y_{in}}{\sum_{j\neq i}^{n}y_{ij}}\bigg{)}.$
The data generating process above is based on the standard relationship between the Dirichlet distribution and a collection of independent gamma variables: if $y_{ij}\sim Gamma(\alpha_{ij},1)$ for $i=1,\ldots,n$ and $j=1,\ldots,i-1,i+1,\ldots,n$, then the compositions
$\mathbf{x}_{i}^{*}=\bigg{(}\frac{y_{i1}}{\sum_{j\neq i}^{n}y_{ij}},\frac{y_{i2}}{\sum_{j\neq i}^{n}y_{ij}},\ldots,\frac{y_{i(i-1)}}{\sum_{j\neq i}^{n}y_{ij}},\frac{y_{i(i+1)}}{\sum_{j\neq i}^{n}y_{ij}},\ldots,\frac{y_{in}}{\sum_{j\neq i}^{n}y_{ij}}\bigg{)}\sim Dir(\alpha_{i1},\alpha_{i2},\ldots,\alpha_{i(i-1)},\alpha_{i(i+1)},\ldots,\alpha_{in}).$
We consider networks of sizes $n=\{30,50,70,100\}$ with $K=3$ clusters and low and medium level of parameter homogeneity, as defined in Section 5 and Appendix D. To better understand how well the model performs when some proportion of the edges is zero-weighted, we test scenarios with $p_{0}=\{0,0.01,0.05,0.1,0.25\}$. Similar to the simulation studies in Section 5.2 and 5.3, we use the Frobenius distance between the true Dirichlet concentration parameter matrix $\mathbf{A}$ and its estimate $\hat{\mathbf{A}}$ to assess the parameter estimation performance, and ARI between the true and the estimated partition for clustering performance.
The results are provided in Figure 4 and Figure 5 respectively, where each boxplot is based on 50 synthetic networks. In Figure 4, we observe that as the proportion of zero-weighted edges increases, the accuracy of parameter estimates deteriorates on average. This result is expected since, when a higher percentage of zeros is present, fewer informative non-zero weights contribute to the parameter estimation, leading to reduced accuracy. Additionally, we note that the impact of zero-weighted edges on parameter estimation is more moderate in cases with low homogeneity compared to cases with medium homogeneity.
(a) Low homogeneity
(b) Medium homogeneity
Figure 4: Assessment of parameter estimation performance of DirSBM using the Frobenius distance between the true parameter matrix $\mathbf{A}$ and its estimate $\mathbf{\hat{A}}$ for networks with varying percentage of zero-weighted edges, based on 50 artificial data sets with $K=3$ and 4(a) low and 4(b) medium level of Dirichlet parameter homogeneity.
(a) Low homogeneity
(b) Medium homogeneity
Figure 5: Assessment of clustering performance of DirSBM using the adjusted Rand index (ARI) between the true and the estimated partition for networks with varying percentage of zero-weighted edges, based on 50 artificial data sets with $K=3$ and 5(a) low and 5(b) medium level of Dirichlet parameter homogeneity.
Referring to Figure 5, we observe that there is a simultaneous reduction in the quality of the clustering structure recovery as the percentage of zero-weighted edges increases, which is expected as poorer parameter estimation makes it more challenging to correctly partition the nodes. Nonetheless, the reduction is quite gradual, and the performance with a small percentage of zero-weighted edges is close to that with no zero-weighted edges (see also Section 5.3), in particular in larger networks. The clustering performance of the model appears to be robust to increasing percentage of zero-weighted edges in larger networks with low parameter homogeneity. It is also worth noting that the zero-weighted edges are drawn uniformly across all pairs of cluster labels, meaning that they do not inform the partition as absent edges might in other stochastic block models.
@@ 6.1 Erasmus programme data
The Erasmus programme is the most popular university student exchange programme in Europe, spanning thousands of Higher Education institutions in more than 30 countries. Gadár et al., (2020) have produced an extensive data set of student, staff and teacher mobility in the Erasmus network by combining data from a number of different sources. The aim of our analysis is to understand global mobility patterns in the Erasmus network, focusing on student exchanges. Due to significant differences in the student populations of the European countries, analysis of raw counts of students per pair of countries may not provide particular insights regarding the popularity of destinations, as the biggest countries dominate simply due to their overall hosting capacity.
After aggregating the student numbers to country level for the 2012-2013 academic year, a directed network with countries as nodes and proportions of students as edge weights is constructed, with the proportions summing to 1 for the sending countries. The Erasmus network is very dense ($\approx 87\%$ of all possible edges are present), and there is no reason to believe that the remaining edges are not allowed to exist since all countries in the Erasmus network are allowed to exchange students. Therefore, instead of treating the edges as absent, and as discussed in Section 2.3, we assign a value of $0.001$ to the interactions with zero count and treat the network as fully connected.
@@
Figure 6(a) illustrates the solution with 3 clusters on a map of Europe, noting the membership of each cluster on the side as smaller countries are difficult to see on the plot. Figure 6(b) represents the total percentage flows between the clusters, based on the estimated matrix $\mathbf{\hat{V}}$. The estimate of the parameter matrix $\mathbf{A}$ as well as the expected node-to-node and cluster-to-cluster exchange matrices $\mathbf{\hat{W}}$ and $\mathbf{\hat{V}}$ from Equations (5) and (6) respectively are provided in E.
Figure 6: Clustering solution for DirSBM with 3 clusters: 6(a) map of Europe with countries coloured by cluster assignment; 6(b) chord diagram of total percentage flows between clusters of countries based on $\mathbf{\hat{V}}$. On the left, S1, S2 and S3 denote the sending cluster (1, 2 and 3). Percentages departing from the sender sum up to 100. On the right, R1, R2 and R3 denote the receiving cluster (1, 2 and 3).
@@
Figure 7: Clustering solution for DirSBM with 5 clusters: 7(a) map of Europe with countries coloured by cluster assignment; 7(b) chord diagram of total percentage flows between clusters of countries based on $\mathbf{\hat{V}}$. On the left, S1-S5 denote the sending cluster (numbered 1 to 5). Percentages departing from the sender sum up to 100. On the right, R1-R5 denote the receiving cluster (numbered 1 to 5).
@@
The second best solution according to ICL with $K=5$ is presented in Figure 7, and the parameter estimates as well as matrices $\mathbf{\hat{V}}$ and $\mathbf{\hat{W}}$ are available in E. The red cluster, now consisting of only Germany, Spain and France, is the cluster of countries receiving the highest percentages of students from the countries of all clusters, ranging between the expected 8.1% to 19.6% per country pair, and the second largest total shares, dominated by the blue cluster by a small margin. It should be noted that the total shares received by the red cluster are only split between 3 countries instead of 11 like in the case of the blue cluster. This cluster is also characterised by retaining 39.1% of its students, and sending 43.3% to the blue cluster, implying a lack of diversity of exchange destinations among its students. The blue cluster can be described as containing reasonably popular exchange destinations, but not quite as popular as the red countries, and, similarly to the red cluster, the majority of students go on exchange within the cluster itself (42.2%) or to the red cluster (38.1%). The green cluster, despite consisting of 9 countries (8 Eastern European countries and Cyprus), does not attract large shares of students (total cluster exchange shares are between 2.4% and 10.6%), which may be due to the fact that such countries are less well known exchange destinations. Green countries have a somewhat more diverse preference profile in comparison to the first two clusters, as the parameter estimates are much closer to each other, but there is still a strong preference for the red and blue cluster countries. Purple countries, which are Switzerland, Hungary, Ireland, Norway and Turkey, can be described as medium-preference exchange destinations. They have a strong preference for the red and blue cluster countries, and very limited interest in countries within its own cluster, as well as in green and orange clusters. The orange cluster is the one having the most homogeneous preferences, as indicated by the smallest differences between the parameter estimates, and its countries tend to be among the least popular destinations for exchange. This can possibly be explained by the fact that these countries are either very small, are located very far from the centre of Europe or are not members of the European Union (as of the 2012-2013 academic year). The green, purple and orange clusters retain very small shares or transfers, 14.2%, 2.4% and 5.6% respectively, while the majority of students from these clusters tend to go on exchange to countries in the red and blue clusters. A final observation is that there seems to be a connection between the exchange preferences of Erasmus students and the level of economic development of the hosting country or the prestige of going on exchange to specific destinations. The countries with stronger economies tend to be grouped together and to exhibit higher expected country-to-country exchange proportions, whereas countries with weaker economies or newer members of the programme, such as Turkey and Croatia, tend to be separated from the more established members.
@@ 6.2 London bike sharing data
We illustrate DirSBM in application to the London bike sharing data from 2014 provided by the Transport for London, (2016). We start by counting the number of trips taken between pairs of stations and identifying the busiest stations for our analysis. To do so, we look at the top 100 start and end stations in terms of volume of bikes exchanged and consider their intersection, i.e. the stations that are both among the most popular start and end points of the journey, resulting in an overall network of 85 stations. The trips only between these 85 stations account for over 10% of the total year’s trips. The network is very dense, with just under 2% of edges having zero weights, hence, similarly to the Erasmus programme network from Section 6.1 and as discussed in Section 2.3, we set the value of the zero-weighted edges equal to $0.001$. We then compute the compositional edge weights by dividing the count weights by the total number of trips taken from the start station.
We fit the DirSBMs with number of clusters between 1 and 8 clusters and select the optimal number of clusters based on the ICL. According to the criterion, the solution with $K=4$ is optimal (ICL$=27866$), followed by the solution with $K=3$ (ICL$=27608$). Figure 8 illustrates the solution with 4 clusters on a map of central London (an interactive version of the map is available here) as well as the chord diagram for total exchange proportions between clusters. The Dirichlet parameter matrix estimate $\mathbf{\hat{A}}$ as well as matrices $\mathbf{\hat{V}}$ and $\mathbf{\hat{W}}$ are provided in F.
@@
Figure 8: Clustering solution for DirSBM with 4 clusters, as selected by ICL, on London bike sharing data: 8(a) map of Central London with bike stations coloured by cluster assignment. Image created using Google My Maps software, with the interactive map available here; 8(b) chord diagram of total percentage flows between clusters of stations based on $\mathbf{\hat{V}}$. On the left, S1-S4 denote the sending cluster (numbered 1 to 4). Percentages departing from the sender sum up to 100. On the right, R1-R4 denote the receiving cluster (numbered 1 to 4).
@@ 7 Discussion
In this paper, we introduced DirSBM, an extension of the stochastic block model to directed weighted networks with compositional edge weights that is based on direct modelling of compositional weight vectors using the Dirichlet distribution. To the best of our knowledge, no counterparts currently exist in the literature, making the proposed clustering methodology a novel contribution. We have developed an inferential procedure based on a variant of the classification expectation-maximisation algorithm for hybrid likelihood. Model selection is addressed using the integrated completed likelihood criterion. We have also proposed an alternative framework for clustering for composition-weighted networks via a log-ratio transformation of the data and subsequent application of an existing weighted SBM to the transformed data. Through simulation studies, we have shown the effectiveness of DirSBM in terms of parameter estimation and recovery of the clustering structure in a variety of settings, and have explored the use of ICL for model selection, which has proven to work reasonably well. The code implementing the modelling framework in R can be found on GitHub.
The modelling approach can be extended in multiple directions. As mentioned at the start of Section 2, the proposed DirSBM models the network’s latent structure based on the sending proportions of the nodes, meaning that the resulting groups are associated with the structure of the nodes acting as senders. However, the latent structure based on the receiving proportions could be quite different as often the nodes behave differently depending on the role they play in the interaction, as noted in Airoldi et al., (2008). Exploring the clustering based on receiving proportions or combining sending and receiving information by extending DirSBM to a two-layer network, with one layer for sending and another for receiving proportions, would be an interesting avenue for future research.
Another important extension to the DirSBM is introducing structural zeros, to enable the ability to model the absence of guaranteed edges in the network. One could use the idea from Tsagris and Stewart, (2018) that concerns Dirichlet regression with zeros, where indicators are used to “discard” zero entries of compositional vectors and subsequently fit the model to the non-zero subset of compositional data. As well as making the model more realistic and applicable to a wider range of real world data, such an extension could lead to numerically more stable performance, since the dimension of the Dirichlet observations would not scale with the number of nodes.
@@
Other developments could include investigating more educated and efficient initialisation strategies for the proposed algorithm. Although random initialisation allows to explore the solution space rather well by considering the most diverse set of starting points, thus reducing the risk of convergence to a local maximum, it does so at a computational cost. This cost could potentially be reduced if the algorithm was to be run from fewer, or ideally a single, more informed initial clustering allocation set. Additionally, alternative estimation methods could also be explored in an attempt to reduce the run times. For instance, the profile pseudo-likelihood approach as proposed by Wang et al., (2023) might offer computational advantages. Unlike Marino and Pandolfi, (2022), where a classification step is implemented within each iteration of the algorithm, in Wang et al., (2023) the columns of the data matrix are classified only after the EM has converged. While such an approach could enhance speed for most SBMs with conditionally independent edges or edge weights, its efficacy for DirSBM is less certain due to the dependence of individual edge weight vectors on the global network latent structure. Any implementations aiming to significantly reduce the computational burden would need to take into account this characteristic of DirSBM and possibly target the estimation of the Dirichlet concentration parameters. Nonetheless, further investigations in this direction could prove fruitful.
Lastly, other distributions for compositional data could be considered to model the sets of proportional edge weights in the network, such as the generalised Dirichlet distribution, which could be a less restrictive approach to capturing compositional variance (Connor and Mosimann,, 1969).
Acknowledgements
This work was supported by the Insight Research Ireland Centre for Data Analytics (Grant number 12/RC/2289_P2).